# Fine-grained Hallucination Detection and Editing for Language Models

### 17 Jan 2024 v2

@@ 4.1 Human Annotation on Hallucinations
To evaluate LMs’ factuality in the responses to diverse queries, we collect 150 information-seeking queries, spanning three different categories. First, we sample 50 knowledge-intensive queries from the Open Assistant Köpf et al. (2023) dataset, by prompting GPT-4 to judge whether each query from the dataset requires world knowledge or not. We, the authors further compose 50 instruction-following queries that require more complex reasoning and knowledge. Lastly, to enable more controlled evaluations and analysis, we synthetically create 50 queries that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template. See examples in each category in Appendix Table 6.
Annotation details.
We prompt ChatGPT Ouyang et al. (2022) and Llama2-Chat 70B Touvron et al. (2023) in a zero-shot manner and collect 300 responses to our diverse information-seeking queries. We recruited 15 students (five undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Due to the complexity of the task, we conducted 45-minute in-person and virtual training sessions with them. Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation. For each LM response, we assign two annotators. We present model-generated responses, input information-seeking queries, and the top 10 web search documents returned by Google Search given the original input queries. We ask human annotators to first mark a span, annotate a category of error, and suggest an edit to fix the errors for the  Entity  and  Relation  types. Our annotation interface and more details of annotations are available in Appendix A.
@@
Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 60.8% and 74.7% of the responses of ChatGPT and Llama2-Chat include at least one factual hallucination, respectively.  Entity  is the most widely recognized error type, making up 46.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements which make up 16.8% and 15.4% of the errors detected respectively. We found that the error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat 70B. We observe fewer errors in Open Assistant subsets, possibly because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How tall is the Burj Khalifa?), which is often memorized by powerful LMs Mallen et al. (2022). On the other hand, the newly annotated instructions and WebNLG data show a higher percentage of errors.
Figure 4: Overview of data creation process in Fava.
Our annotators also agreed that our taxonomy covers most of the error types in LM responses. Furthermore, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in error detection across passages, with 75.1% agreement in error detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages. Appendix Table 8 shows examples of our annotations.
@@ 5 Model: Fava
Our benchmark analysis reveals the urgent need to study and detect such diverse types of hallucinations.
@@
We introduce a new retrieval-augmented LM, Fava (FAct Vericaton with Augmentation). Fava is trained on high-quality synthetic training data, and at inference, it identifies and fixes fine-grained hallucinations, incorporating retrieved knowledge.
5.1 Overview
Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and if possible edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1.
@@ 5.1 Overview
While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2). Reliance on black-box proprietary API models also hurt reproducible. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.2) and fine-tune a powerful expert LM consisting of 7 billion parameters (Section 5.3).
5.2 Training Data Creation
@@
To train our $\mathcal{M}_{edit}$ model, we require a large number of erroneous LM outputs $y$ paired with edited versions $y^{*}$ with fine-grained error types and edits. Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al. (2022); Wang et al. (2023b); Asai et al. (2023), we introduce a new data creation method grounded on our fine-grained hallucination taxonomies. We prompt GPT-4 Achiam et al. (2023) and ChatGPT to diversify and noise a Wikipedia passage by inserting different types of hallucinations.
In particular, our data creation pipeline consists of three steps: seed passage generations, error insertions, and postprocessing. The first two processes create an input sequence with perturbed diverse error annotations, and postprocessing remaps those special tokens to generate the erroneous output $y$ and output with fine-grained detection results $y^{*}$. We use the original Wikipedia passage as well as passages retrieved by an off-the-shelf system as $\mathbf{C}$. Figure 4 shows the overview of the pipeline.
Seed passage generation.
@@
We randomly sample a Wikipedia article $c$ and use it as our gold reference passage. To produce model-generated seed passages, we diversify the sampled article by paraphrasing and transferring the text to another genre via an LM. Then, we prompt ChatGPT to create a new genre of the text $t$, specifying to include information only found in the Wikipedia article $c$ to maintain factuality. We randomly sample one of the pre-specified genre types (e.g. blog article, question answer pair, tweet, etc.) for each passage. To further diversity in the dataset, we also add examples of question answering (QA) data. In particular, we prompt ChatGPT with two demonstrations from the NaturalQuestions Kwiatkowski et al. (2019) dataset to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles. See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 9.
Error insertion.
@@
In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types. Therefore, we use ChatGPT and GPT-4 interchangeably for six different types. [^3] We use ChatGPT for the Type 1 and Type 3 while using GPT-4 for other types and insert factual errors one by one into our diversified text. In particular, given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retraining previously inserted errors and original text. The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution.
After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
Post processing.
@@
We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. At this stage, we also filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types). As a result, we have a training instance $(c,y,y^{*})$, consisting of the gold context (original Wikipedia paragraph) $c$, erroneous LM output $y$, and output with error tags and correct editing $y^{*}$. We also retrieve additional four relevant paragraphs from Wikipedia using Contriever Izacard et al. (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$. As this process does not involve human annotations, we can easily scale up training instances under an API cost budget.
Statistics of data.
@@
We generated a total of 35,074 training instances, 30,0074 of which are based on Wikipedia passages, and 5,000 are based on QA pairs described above. All of the error types are almost equally distributed (roughly 15% each). We present the error distributions in Appendix Table 10. We conduct a detailed human evaluation in Section 7.2. On average, 3.1 errors are inserted for the Wikipedia passage-based subsets, and 1.4 errors are inserted for the QA pairs.
@@ 5.3 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$.
At inference time, we retrieve the top five documents from Wikipedia, [^4] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder. using Contriever-MSMARCO Izacard et al. (2022), and insert them together with an LM output that may include factual errors. The model identifies factual errors, marks phrases or sentences for deletion, and suggests edits for improving the factuality of the text.
@@ 6 Experiments
We evaluate Fava and state-of-the-art LMs with and without retrieval on the fine-grained hallucination detection (Section 6.1) and editing tasks (Section 6.2). We further conduct human evaluations on our best models (Section 6.3).
ChatGPT Generations Llama2-Chat 70B Generations
Editor  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
Rt+CGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
Fava (ours) 43.1 25.0 57.1 16.7 70.5 35.3 46.5 78.2 56.2 34.5 27.7 34.5 67.0 43.4 46.5 80.6
Table 2: Overall fine-grained detection performance. CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
@@ 6.1 Experiments for Hallucination Detection
Evaluation data.
We use our new benchmark, consisting of 364 annotated passages, [^5] In total we collected 471 annotated passages including our experimental batches, and we excluded earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.3.
Baselines.
For fine-grained hallucination detection tasks, we test three baselines. ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. Rt-ChatGPT uses the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^6] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
6.2 Experiments for Editing
Evaluation data and function.
For editing evaluations, we use two open-ended text generation data and evaluate models’ factuality based on two metrics designed to measure the factuality of text. FActScore Min et al. (2023) prompts ChatGPT and GPT 3-davinci 003 to decompose a response into a set of atomic facts and verify factuality for each of them using passages from a designated Wikipedia article. We use the original biography generation task proposed with FActScore for evaluation.
Baselines
For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively). As in the fine-grained error detection task, we prompt the models with carefully designed prompts and two demonstrations. For all baselines, we retrieve five documents consisting of one document retrieved by string matching based on the entity name and four documents retrieved by Contriever, which are reranked at test time, using a small cross-encoder. [^7] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2
@@ 6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect annotated data and editing errors due to the potential subjectivity of annotations and may be affected by the performance of factuality evaluation metrics. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level. Due to the cost of evaluating the factuality of all verification-worthy statements Min et al. (2023), we do not evaluate recall.
@@ 7.1 Results
Detection results.
Table 2 shows the fine-grained detection accuracy of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 12 and 13, respectively. We also provide binary prediction performance.
Table 2 shows that while ChatGPT-based baselines show relatively good binary predictions of existence of factual errors, their fine-grained error type detection performance is significantly lower. Fava shows its strong capabilities of conducting fine-grained error detection, resulting in significant performance improvements from ChatGPT or Ret-ChatGPT on both fine-grained error detection and binary error detection. Specifically, Fava shows high accuracy on error types such as  Contradictory ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable are still limited. Those two error types often require intensive search over many web documents beyond top few passages, while Fava by default only considers top five documents. Our inter-annotator agreements also show that there are more disagreements in those two categories, which might also deflate the performance.
ChatGPT Al-13B Al-7B
@@
Table 3: Overall results on FActScore. Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
Editing results.
Table 3 shows the results on FactScore. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model size. Retrieval-augmented ChatGPT shows small improvements by editing Alpaca outputs, editing its output with retrieved documents hurts the FActScore. We found that ChatGPT is often confused with different error types and exhibits tendencies of over-predicting contradictory or subjective statements. Among other models, our methods show the largest gains in terms of factuality, showing the effectiveness of editing, in addition to its strong capability of providing fine-grained error types.
Human evaluation results.
Our human evaluation results are shown in Table 4. As in automatic evaluations, our method shows significantly better performance than Rt+ChatGPT on both editing and detection and recognizes more errors than retrieval-augmented ChatGPT. These results further demonstrate the strong capabilities of Fava detecting factual errors in LM outputs.
Model avg. $|\mathbf{E}|$ Detect (%) Edit (%)
Rt+ChatGPT 1.9 23.9 17.2
Fava (ours) 2.4 55.7 63.5

### 21 Feb 2024 v3

@@ 4 Benchmark: FavaBench
The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 4.
  • Knowledge-intensive queries sampled from the Open Assistant dataset Köpf et al. (2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
  • A sample of Open QA prompts from the No Robots dataset Rajani et al. (2023).
  • Instruction-following prompts requiring more reasoning and knowledge curated by the authors.
  • Synthetically-created prompts 50 prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template.
Annotation details and quality.
We obtain responses to the collected source prompts using ChatGPT (gpt-3.5-turbo-0301;  Ouyang et al. 2022) Llama2-Chat 7B and Llama2-Chat 70B Touvron et al. (2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts. We recruited 20 students (ten undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Each instance is annotated by two annotators who completed 45-minute in-person and virtual training sessions. [^6] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation. Our annotation interface and details are in Appendix A. To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages. Appendix Table 6 shows examples of our annotation results.
@@
Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 59.8%, 70.2% and 64.9% of the responses of ChatGPT, Llama2-Chat 7B and Llama2-Chat 70B include at least one hallucination, respectively.  Entity  is the most widely recognized error type, making up 48.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements, which make up 14.1% and 18.1% of the detected errors, respectively. The error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat models. Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs Mallen et al. (2022).
Figure 4: Overview of high-quality synthetic data generation process in Fava. Fava leverages powerful instruction-tuned models to carefully insert errors into factually accurate statements and produces diverse error types based on our proposed taxonomy.
@@ 5 Model: Fava
We introduce Fava (FAct Vericaton with Augmentation), a model for fine-grained hallucinations detections and editing. Fava is trained on high-quality synthetic training data to identify hallucinations, incorporating retrieved knowledge.
@@
Overview. Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and, if possible, edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1. While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2). Reliance on such API models also hurt reproducibility. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful expert LM on the generated data (Section 5.2).
5.1 Synthetic Training Data Curation
@@
To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context (original Wikipedia paragraph), $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$. Fava leans to take $(c,y)$ and generate $y^{*}$. Yet, such data is not publicly available, and our annotation is expensive to scale up. Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al. (2022); Wang et al. (2023b); Asai et al. (2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies.
Specifically, our data creation pipeline consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements informed by our our hallucination taxonomy to generate erronous passages $y$, and Step 3: post-processing which prunes and curate the final training examples. Figure 4 shows the overview of the pipeline.
Seed passage generation. The goal in this step is to collect and generate a large, diverse range of seed passages to ensure Fava can handle a wide range of text types while maintaining control over factuality. Later on, errors are inserted into these seed passages to generate erroneous text, $y$. We base our seed passage collection on Wikipedia articles and available Question Answering (QA) datasets Kwiatkowski et al. (2019) since they are both good sources to evaluate factuality. However, the text genre is quite limited. To diversify the passage genres, we paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats.
@@
More specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages. We then randomly sample one of the pre-specified genre types (e.g. blog article, tweet, etc.) for each passage and use ChatGPT to paraphrase the original passage into the given genre. In addition, we use 5,000 examples from QA datasets to further diversify the passage collections. In particular, we prompt ChatGPT with two QA demonstrations to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles. See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 7.
@@
Error insertion. The goal of this step is to simulate erroneous LM generation $y$, which are passages with factual errors. To control error distributions in the training data and incorporate our hallucination taxonomy, we prompt LMs to insert errors from our taxonomy one by one. In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types. Therefore, we use ChatGPT and GPT-4 interchangeably for six different types. [^7] We use ChatGPT (gpt-3.5-turbo-0301) for the Type 1 and Type 3 while using GPT-4 for other types. Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors and original text. The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution. After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
@@
Post-processing. We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. We filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types). We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO Izacard et al. (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
@@
Statistics of data.  After analyzing ablations with varying training instance sizes (Section 7.2), we generated 35,074 training instances,  [^8] While we see performance improvements up to 30k, the API costs for training data generation were already \$3k, reaching our initial budget. 30,0074 based on Wikipedia passages, and 5,000 based on QA pairs. All of the error types are almost equally distributed ($\sim$15 % each) with 3.1 errors inserted for Wikipedia subsets, and 1.4 errors inserted for the QA pairs on average. We present the error distributions in Appendix Table 8 and a detailed human evaluation in Appendix C.4.
@@ 5.2 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using the standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$. We empirically found initializing Fava with this model gives slightly better performance than the pre-trained Llama2.
At inference time, we retrieve the top five documents from Wikipedia, [^9] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder. using Contriever-MSMARCO, and input them together with an LM output that may include hallucinations. The model identifies hallucinations, marks phrases or sentences for deletion, and suggests edits.
@@ 6 Experiments
Generator: ChatGPT Generator: Llama2-Chat 70B
Editor  ent  rel  con  inv  subj  unv OA Bi  ent  rel  con  inv  subj  unv OA Bi
ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 60.8 55.5 60.0 21.2 15.4 2.0 25.0 42.5 74.2
Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 79.6 57.3 34.5 27.7 52.2 31.25 43.4 47.2 80.3
Table 2: Fine-grained detection F1. OA and Bi indicates overall and binary predictions.
@@ 6.1 Experiments for Hallucination Detection
Evaluation data.  We use our new benchmark, consisting of 902 annotated passages, [^10] We collected 1010 annotated passages including our experimental batches, excluding earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.2.
Baselines.  For fine-grained hallucination detection tasks, we test three baselines. ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. GPT4 (gpt-4) does the exact same but prompts GPT4 instead.We were unable to include GPT4 with retrieval in our baseline due to high costs. Since Fava uses retrieval augmentation, for comparability we include Rt+ChatGPT. This baseline prompts ChatGPT using the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^11] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
6.2 Experiments for Hallucination Editing
Evaluation data.  For editing evaluations, we use open-ended text generation data and evaluate models’ factuality based on a metric designed to measure the factuality of text, FActScore Min et al. (2023). This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article. We use the original biography generation task proposed with FActScore.
Baselines.   For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively), and prompt them with carefully designed prompts and two demonstrations. For all baselines, we retrieve five documents consisting of one document retrieved by string matching based on the entity name and four documents retrieved by Contriever, which are reranked at test time, using a small cross-encoder. [^12] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2
@@ 6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect and refine hallucinations, due to the potential subjectivity of our annotations or the performance of the factuality evaluation metric. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level. Due to the cost of evaluating the factuality of all verification-worthy statements Min et al. (2023), we do not evaluate recall.
@@ 7.1 Results
Detection results.  Table 2 shows the fine-grained detection accuracy, namely F1 scores, and binary prediction F1 of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 11 and 12, respectively. Fava significantly outperforms ChatGPT, Ret-ChatGPT, or GPT-4 on both fine-grained error detection and binary error detection. Fava shows high accuracy on error types such as   Sentence ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable is still limited. Those two error types often require intensive search over many documents beyond the top few ones, while Fava by default only considers the top five documents.
Generator
Editor ChatGPT Al-13B Al-7B
@@
Table 3: Overall Editing results on FActScore Min et al. (2023). Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
Editing results.  Table 3 shows the results of the editing task on FActScore. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model in size. We found that ChatGPT is often confused with different errors and has a tendency to over-predict contradictory or subjective statements. Fava shows the largest gains in FActScore, showing the effectiveness of editing, in addition to error type detection.
Human evaluation results.  Our human evaluation results are shown in Appendix Table 14. Fava shows significantly better performance than Rt+ChatGPT on both editing and detection scoring 46.3% and 31.8% more than Rt+ChatGPT on each task respectively. Furthermore, Fava recognizes more errors, 2.4 errors per passage, than retrieval-augmented ChatGPT, 1.9 errors per passage. These results further demonstrate the strong capabilities of Fava detecting factual errors in LM outputs.